# Interpreting Key Mechanisms of Factual Recall in Transformer-Based Language Models

### v1

@@ Appendix B Details on Solving the Linear Regression
We present the regression solutions for Eq.7 in GPT-2 small on the country-capital task. The coefficients for other tasks and models will be available in our code repository. Table 5 presents the learned coefficients for each layer, along with the validation set loss in our 4-fold cross-validation. The optimization consistently converges to similar solutions for each layer with minimal loss. Although a few weights in one of the four experiments might diverge from the other three a little, the overall results remain stable enough to support qualitative analysis. Furthermore, it is natural for the validation MSE loss to rise as the layer depth increases, given that the vectors in deeper layers possess larger norms. As MSE is a quadratic loss whose penalty is proportional to the square of the error, to alleviate any concerns regarding the convergence, we also present the Absolute Percentage Error (APE) in Table 6, expressed as $\frac{\sqrt{\text{MSE}(\tilde{\textbf{m}}^{l},\textbf{m}^{l})}}{||\textbf{m}^{l}||}\cdot 100\%$. The APE values for each layer hover around a mere 1%, indicating minimal deviation.
@@
Appendix C Identifying the MLP Layer for “Function Application”
In our main text, for simplicity’s sake, we followed Merullo et al. [2023] to identify the layer where the probability of $Y$ surpasses that of $X$ as the point of function application. However, this principle is not accurate. A more general principle would involve examining the layers “around” the point where the probabilities of $X$ and $Y$ begin to diverge significantly. To illustrate this, Figure 7 showcases the zero-shot probabilities of $X$ and $Y$ across the layers in GPT-2 medium and GPT-2 large models.
Appendix D GPT-2 small on the “Product-Developer” task
Figure 8: We test GPT-2 small on the product-developer task. (a) The probability of $X$ and $Y$ when early unembed each layer’s outputs. The shaded regions indicate the standard deviation. (b) The effect of patching path $\textbf{a}^{l,h}_{t=-1}\rightarrow\textbf{r}^{11}_{\text{post}}$. (c) Value weighted attention pattern of L10H0. (d) L10H0 is a mover head.
Appendix E Experiments on OPT-1.3B
OPT-1.3B [Zhang et al., 2022] consists of 24 layers with 32 heads per layer. In Figure 13, we present the probability dynamics across layers, along with path patching experiments. From Figure 13(a), the argument passing is observed in layer 19. Figure 13(b) reveals that L19H15 in the zero-shot settings acts as the first influential argument passer. While L17H8 and L18H7 also serve as argument passers, their impact on increasing the probability of $X$ is diminished within the layer. Their activation is enhanced in few-shots scenarios, making the probability of $Y$ emerge in the residual stream early, as shown in Figure 15(b) and (c), which is similar to the behavior of L8H11 in GPT-2 small.
Recall that in §3.4, two key factors prompted us to realize that $\textbf{b}^{10}$ directs from $X$ to $Y$ within the MLP10 of GPT-2 small: Firstly, $X$ and $Y$ are equally important tokens that occupy most of the probability distribution. Secondly, after MLP10, the probability of $X$ decreases, but that of $Y$ increases. We have emphasized that these two factors do not always occur, and interpreting the mechanisms of MLPs requires a case-by-case analysis, depending on the model, task, or even prompts.
Token Node $\textbf{r}^{20}_{\text{mid}}$ $\textbf{r}^{20}_{\text{post}}$
not 1.32% 4.97%
London 4.17% 3.21%
England 67.40% 74.40%
Table 7: Probability of certain tokens decoded from several nodes in OPT-1.3B.
Regarding the final layer, as illustrated in Figure 15, OPT-1.3B also demonstrates anti-overconfidence tendencies, suppressing accurate predictions across zero-shot, one-shot, and two-shot scenarios. By employing the anti-suppression techniques outlined in §4, we effectively alleviate the anti-overconfidence mechanism employed by OPT.
Appendix F Future Works
This section aims to outline several research questions that remain unexplored. We present hypotheses for each and leave them for future research endeavours.
Firstly, in §3.2, we directly start with that task-specific heads are activated by task semantics. However, we omit to explore how the model constructs this task semantics within its shallow layers. We attempted to trace the information flow from the query and key vectors of the mover heads, only to find that the heads influencing these vectors are scattered, and the individual impact of a head is minimal. We posit that the formation of the task semantic occurs through collaborative efforts across multiple circuit paths. Thus, attempting to patch a single pathway might be futile, as another pathway could compensate. Collaborative circuit paths have long posed a challenge for circuit discovery methods Conmy et al. [2023]. Additionally, this seems related to the Hydra effect in neural networks McGrath et al. [2023].
Secondly, in this paper, we did not study the origins of these mechanisms. Take, for instance, the behavior of an MLP that redirects the residual stream to its expected answer. This phenomenon seems to correlate with the residual connections within the Transformer architecture. When a module handles an input, denoted as $x$, to attain a target, $y$, the most effective strategy entails minimizing loss by structuring the output as $y-x$ instead of aligning the output with the direction of $y$. This manner is more effective despite the potential for outputs to seem orthogonal or display a negative cosine similarity with $y$. Furthermore, the anti-overconfidence mechanism of the final layer is intriguing. We hypothesize that this layer acts as a final “defense line,” employed by the model to mitigate overconfidence. A confident prediction could result in a significant loss if it outputs an incorrect format or enters the wrong answer domain.
We also notice that the direction of the intercept b in our paper encodes the semantics of the prediction target. This observation resonates with findings in various studies where modifying the outputs of MLPs using activations from other prompts resulted in corresponding changes in model outputs [Zhang et al., 2024, Merullo et al., 2023, Todd et al., 2023, Hendel et al., 2023]. Therefore, our analysis method for decomposing MLP outputs has the potential to prove beneficial across various applications requiring controllability, such as role-playing [Tu et al., 2024], stylized dialogue [Li et al., 2021], and model jailbreak [Zou et al., 2023], which we plan to study in future works.

### v4

@@ Appendix I Details on solving the linear regression
We present the regression solutions for Eq.6 in GPT-2 small on the country-capital task. Table 7 presents the learned coefficients for each layer, along with the validation set loss in our 4-fold cross-validation. The optimization consistently converges to similar solutions for each layer with minimal loss. Although a few weights in one of the four experiments might diverge from the other three a little, the overall results remain stable enough to support qualitative analysis. Furthermore, it is natural for the validation MSE loss to rise as the layer depth increases, given that the vectors in deeper layers possess larger norms. As MSE is a quadratic loss whose penalty is proportional to the square of the error, to alleviate any concerns regarding the convergence, we also present the Absolute Percentage Error (APE) in Table 8, expressed as $\frac{\sqrt{\text{MSE}(\tilde{\textbf{m}}^{l},\textbf{m}^{l})}}{||\textbf{m}^{l}||}\cdot 100\%$. The APE values for each layer hover around a mere 1%, indicating minimal deviation.
@@
Appendix J Limitations and future works
Our paper left several research questions unexplored. We present hypotheses for each and leave them for future research endeavours.
Firstly, in Section 3.2, we directly start with that task-specific heads are activated by task semantics. However, we omit to explore how the model constructs this task semantics within its shallow layers. We attempted to trace the information flow from the query and key vectors of the mover heads, only to find that the heads influencing these vectors are scattered, and the individual impact of a head is minimal. We posit that the formation of the task semantic occurs through collaborative efforts across multiple circuit paths. Thus, attempting to patch a single pathway might be futile, as another pathway could compensate. Collaborative circuit paths have long posed a challenge for circuit discovery methods [3]. It might be related to the hydra effect in neural networks [19]. Additionally, the mechanism of in-context learning boosting model confidence also needs future investigation.
Secondly, the origins of these mechanisms need more research endeavors in the future. Take, for instance, the behavior of an MLP that redirects the residual stream to its expected answer. This phenomenon seems to correlate with the residual connections within the Transformer architecture. When a module handles an input, denoted as $x$, to attain a target, $y$, the most effective strategy entails minimizing loss by structuring the output as $y-x$ instead of aligning the output with the direction of $y$. This manner is more effective despite the potential for outputs to seem orthogonal or display a negative cosine similarity with $y$.
Appendix K Broader impacts and safety issues
This paper reveals many intriguing phenomenons and a novel analysis method to understand MLPs in language models, which have the potential to enhance language models during inference, design new foundational architecture, and improve training paradigms towards a higher level of intelligence. Moreover, this interpretative endeavour has the potential to steer language models towards alignment with human values, thereby mitigating safety concerns associated with existing large-scale language models. Nonetheless, while striving to understand the black-box model fully, there exists a risk that humans may develop models that intentionally violate human values. This safety issue needs careful consideration.